OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. 15 = 4a - 5 Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. 15 = 4a - 5 @Mathematics

OpenStudy (asnaseer):

ok - how would you solve this?

OpenStudy (anonymous):

r u asking me?

OpenStudy (asnaseer):

yes :-)

OpenStudy (anonymous):

Ummm... wouldn't you take 4-5,then that answer +15?

OpenStudy (asnaseer):

not quite, this is how you should do it: 15 = 4a - 5 [next add 5 to both sides in order to remove the -5 from the RHS] 15 + 5 = 4a - 5 + 5 20 = 4a [next divide both sides by 4] 20/4 = 4a/4 5 = a \lastly swap the LHS and RHS to leave equation in standard form] a = 5 Think of each variable ("a" in this case) as a different kind of fruit. so think of "a" as apples and think of the constant numbers as oranges. you need to collect all fruits of the same type and you cannot merge different fruits (i.e. apples and oranges cannot be merged - or in this case terms in "a" cannot be merged with the constant terms)

OpenStudy (anonymous):

Okay i think i get it now thanky you so much for your help

OpenStudy (asnaseer):

no problems - glad to be of service! :-)